### v1

# Operator algebra of the information-disturbance tradeoff in quantum measurements

Abstract
Quantum measurements can be described by operators that assign conditional probabilities to different outcomes while also describing unavoidable physical changes to the system. Here, we point out that operators describing information gain at minimal disturbance can be expanded into a set of unitary operators representing experimentally distinguishable patterns of disturbance. The observable statistics of disturbance defines a tight upper bound on the information gain of the measurement.
I Introduction
It is well known that there is an information-disturbance tradeoff in quantum mechanics. Any information gain due to a measurement of the system also implies a disturbance of that system [1, 2]. In terms of operators, a quantum measurement necessarily changes states that are not eigenstates of the measurement operator. However, the precise evaluation of this tradeoff using concepts of information theory is surprisingly difficult [3, 4, 5, 6, 7]. Existing literature employs tools from quantum information theory such as mutual information, quantum discord, and mean estimation fidelity to bound the tradeoff between information gain and the disturbance introduced to the system during measurement [9, 10, 11]. These tools tend to focus on the information content of quantum states, making it somewhat difficult to identify the role of the measurement process itself in the definition of the tradeoff. Here, we propose to shift the focus to the operators that describe the measurement process, introducing an analysis of the operator algebra that defines the information-disturbance tradeoff.
In general, quantum measurements can be described as trace preserving linear maps characterized by specific sets of Kraus operators $\hat{M}_{m}$ representing the different measurement outcomes $m$ [12]. Information is gained because the probability of $m$ depends on the input state. This dependence is described by the eigenstates $\{\left|a\right\rangle\}$ of the self-adjoint operators $\hat{M}_{m}^{\dagger}\hat{M}_{m}$. In this work, we take advantage of the fact that the corresponding eigenvalues correspond to the classical conditional probabilities $p(m|a)$ that describe the information gain with respect to the unknown variable $a$. Although full quantum-classical correspondence is only achieved when information is encoded in the basis $\{\left|a\right\rangle\}$, it is worth noting that the measurement process itself defines this basis as the one about which maximal information can be obtained.
The minimal disturbance caused by information gain is described by the changes in quantum amplitudes of the components $\left|a\right\rangle$ associated with the Bayesian updates of the corresponding probabilities. A measurement that achieves this minimal disturbance can be represented by positive self-adjoint operators $\hat{M}_{m}$ with eigenvalues of $\sqrt{p(m|a)}$. We then consider the possibility of analyzing the disturbance by representing the measurement as a superposition of unitary operators [13]. We find that the minimally disturbing measurement described by operators $\hat{M}_{m}$ with eigenstates $\{\left|a\right\rangle\}$ can be characterized by expanding $\hat{M}_{m}$ into a set of orthogonal unitary operations, transforming the representation of $\hat{M}_{m}$ into a superposition of different possible changes to a complementary basis $\{\left|b\right\rangle\}$. This allows us to identify the maximal experimentally observable disturbance associated with the measurement process.
The article is structured as follows. In Section II, we introduce the operator representations of minimally disturbing measurements and show how the representation of information gain transforms into the representation of disturbance. In Section III, we consider the experimental characterization of disturbance and information gain using the two complementary basis systems. In Section IV we formulate a tight bound for the information gain derived from the observed disturbance. It is pointed out that this limit can be used to characterize information leaks in quantum channels. Section V concludes the paper.
II Operator expressions of information and disturbance
A minimally disturbing measurement process that distinguishes a set of orthogonal states $\{\left|a\right\rangle\}$ can be represented by a self-adjoint measurement operator $\hat{M}_{m}$, where $m$ represents the outcome of the measurement related to $a$ by the conditional probabilities $p(m|a)$. The spectral decomposition of the measurement operator is given by
$\hat{M}_{m}=\sum_{a=0}^{d-1}\sqrt{p(m|a)}\left|a\right\rangle\thinspace\left\langle a\right|.$ (1)
The projection operators $\left|a\right\rangle\thinspace\left\langle a\right|$ represent the causal relation between the target observable described by the eigenstates $\left|a\right\rangle$ and the measurement outcome $m$. In this sense, the operator representation is biased in favor of the information extraction represented by the eigenvalues $\sqrt{p(m|a)}$. At the same time, the operator necessarily changes the quantum state that it acts upon by modifying the coherences between the eigenstate components $\left|a\right\rangle$. This modification of the coherences between eigenstates describes the necessary disturbance of the state caused by the extraction of information described by the conditional probabilities $p(m|a)$. It follows that we can identify the physics of this disturbance by changing the representation of the operators $\hat{M}_{m}$ to a description that identifies the precise change of the coherences between the eigenstates $\left|a\right\rangle$.
Mathematically, operators form their own vector space, with the product trace of two operators serving as the inner product. It is therefore possible to adapt the representation of operators to the physics of a given problem. Since we are interested in changes of the quantum state, our operator basis should be composed entirely of unitary operators. It is indeed possible to define a complete basis of $d^{2}$ orthogonal unitary operators acting on states in a $d$-dimensional Hilbert space. These operators are known as the Heisenberg-Weyl group and describe combinations of discrete displacements acting on a pair of mutually unbiased basis sets [14, 15, 16, 17]. In the case of the operator $\hat{M}_{m}$, the situation is simplified by the fact that the operator expansion given by Eq.(1) exists in the $d$-dimensional subspace spanned by the projection operators $\{\left|a\right\rangle\thinspace\left\langle a\right|\}$. The corresponding $d$-dimensional set of orthogonal unitary operators $\{\hat{U}(k)\}$ is given by
$\hat{U}(k)=\sum_{a=0}^{d-1}\exp(i\frac{2\pi}{d}ka)\left|a\right\rangle\thinspace\left\langle a\right|.$ (2)
The phase factors ensure that the product traces of operators with different values of $k$ are orthogonal to each other. The orthogonality relation is given by
$\text{Tr}\bigg(\hat{U}(k)^{\dagger}\hat{U}(k^{\prime})\bigg)=d\delta_{k,k^{\prime}},$ (3)
where $d$ is the trace of the identity obtained for $k=k^{\prime}$. We can now represent the measurement operator $\hat{M}_{m}$ by a linear combination of these unitary operations,
$\hat{M}_{m}=\sum_{k=0}^{d-1}C_{k}\hat{U}(k),$ (4)
where the coefficients $C_{k}$ are given by a discrete Fourier transform of the square roots of the conditional probabilities,
$C_{k}=\frac{1}{d}\sum_{a=0}^{d-1}\exp(-i\frac{2\pi}{d}ka)\sqrt{p(m|a)}.$ (5)
This representation of the measurement operator describes a superposition of experimentally distinguishable changes to the physical properties of the quantum system. In the following, we will consider the experimental characterization of the physical changes described by the measurement operator $\hat{M}_{m}$ and relate them to the information extraction described by the conditional probabilities $p(m|a)$.
III Experimental characterization
Experimentally, the unitary operations can be distinguished by their effects on the Fourier basis $\{\left|b\right\rangle\}$ conjugate to $\{\left|a\right\rangle\}$. The effect of the unitaries is then given by a cyclic shift,
$\hat{U}(k)\left|b\right\rangle=\left|b+k\right\rangle,$ (6)
where the complementary basis is related to the original basis $\left|a\right\rangle$ by a discrete Fourier transform,
$\left|b\right\rangle=\frac{1}{\sqrt{d}}\sum_{a}\exp(-i\frac{2\pi}{d}ka)\left|a\right\rangle.$ (7)
The effect of the measurement operator $\hat{M}_{m}$ on an input state $\left|b\right\rangle$ is given by
$\hat{M}_{m}\left|b\right\rangle=\sum_{k}C_{k}\left|b+k\right\rangle.$ (8)
If a measurement of $\{\left|b^{\prime}\right\rangle\}$ is performed on the output, the joint probability of $m$ and $b+k$ is given by
$p(m,b+k|b)=|\left\langle b+k\right|\hat{M}_{m}\left|b\right\rangle|^{2}=|C_{k}|^{2}.$ (9)
An experimental characterization of the disturbance patterns observed in the complementary basis $\left|b\right\rangle$ thus determines the absolute values of the coefficients $C_{k}$.
As stated above, the coefficients $C_{k}$ represent the discrete Fourier transform of the square root of the conditional probabilities $p(m|a)$ over the index $a$ of the eigenstates $\{\left|a\right\rangle\}$. The dependence of the probability of $m$ on the input $a$ thus characterizes a specific pattern of disturbance in the operator algebra expressed in terms of the coefficients $C_{k}$. To better understand the information-disturbance tradeoff, it is important to remember that the conditional probability $p(m|a)$ defines the Bayesian update of the probabilities of $a$ for the outcome $m$. If there is no prior information available (all $a$ are equally likely), the updated probability is given by
$p(a|m)=\frac{p(m|a)}{\sum_{a^{\prime}}p(m|a^{\prime})}.$ (10)
Here, a probability of $p(a|m)=1$ describes a successful measurement of $a$ for the outcome $m$. In general, the maximal value of $p(a|m)$ provides a measure of the information about $a$ obtained from the measurement outcome $m$. With this in mind, we can now formulate the information-disturbance tradeoff for the measurement operator $\hat{M}_{m}$.
IV Information-disturbance tradeoff of a measurement operator
Eq.(5) describes the relation between disturbance and information gain by relating the complex coefficients $C_{k}$ to the conditional probabilities $p(m|a)$. It should be noted that the real values of $\sqrt{p(m|a)}$ restrict minimally disturbing measurements to $C_{d-k}=C_{k}^{*}$. With this restriction in mind, it is possible to invert the relation to find that
$\sqrt{p(m|a)}=\frac{1}{d}\sum_{k}\exp(i\frac{2\pi}{d}ka)C_{k}.$ (11)
Knowledge of the complex values $C_{k}$ would therefore allow us to reconstruct the conditional probabilities $p(m|a)$ for the measurement outcome $m$. However, only the absolute values $|C_{k}|$ can be observed experimentally. It is therefore interesting to consider the information obtained about $p(m|a)$ from the absolute values $|C_{k}|$. We find that the maximal value of $p(m|a)$ among all $a$ is obtained when the phases of all $C_{k}$ line up so that
$\sqrt{p(m|a)}\leq\frac{1}{d}\sum_{k}|C_{k}|.$ (12)
This inequality should now be applied to the relevant probabilities. Eq.(10) describes the probability of successfully identifying a specific input $a$ based on the measurement outcome $m$. It should be noted that the sum over all inputs $a^{\prime}$ in this relation can be observed experimentally as the probability of observing $m$ for an input state $\left|b\right\rangle$,
$p(m|b)$ $=$ $\sum_{k}p(m,b+k|b)$ (13)
$=$ $\frac{1}{d}\sum_{a^{\prime}}p(m|a^{\prime}).$
We can now express Eq.(12) as an upper bound on the probability $p(a|m)$ referring to the successful determination of $a$ from the measurement result $m$. The bound reads
$p(a|m)\leq\frac{1}{d}\left(\sum_{k}\sqrt{p(b+k|b,m)}\right)^{2},$ (14)
where
$p(b+k|b,m)=\frac{p(m,b+k|b)}{p(m|b)}.$ (15)
The distribution of quantitative disturbances $k$ changing the observable property $b$ determines an upper bound on the probability of successfully identifying $a$ from the measurement outcome $m$. The bound is tight since it can be achieved by any set of real and positive disturbance coefficients $C_{k}=|C_{k}|$. The only way to obtain $p(a|m)=1$ is to have an equal distribution of disturbances, $p(b+k|b,m)=1/d$. Even the slightest preference for a specific value of $k$ defines a non-trivial upper limit of $p(a|m)$. On the other hand, the absence of disturbance given by $p(b|b,m)=1$ limits $p(a|m)$ to $1/d$, corresponding to a random guess of $a$.
V Conclusions
The information gain described by minimally disturbing operators $\hat{M}_{m}$ is described by conditional probabilities $p(m|a)$ that completely characterize the operator. It is therefore possible to characterize the information-disturbance trade-off by transforming the spectral decomposition of the operator $\hat{M}_{m}$ into a superposition of unitary operations that represent the unavoidable disturbance. As we have shown above, the results of this transformation provide deeper insights into the relation between information and disturbance as an intrinsic feature of the measurement process. Eq. (14) describes the experimentally observable bound on information gained about $a$ for a specific pattern of disturbance observed in the complementary property $b$. When characterizing a quantum channel, the observed disturbance of the inputs $\{\left|b\right\rangle\}$ can be used to put an upper limit on the possible leakage of information encoded in $\{\left|a\right\rangle\}$. This limit may be useful in a quick assessment of quantum cryptography protocols, providing the likelihoods of identifying eavesdroppers in BB84 type protocols where the information is encoded in two complementary basis systems $\{\left|a\right\rangle\}$ and $\{\left|b\right\rangle\}$ [18].
It should be emphasized that the complementary basis $\{\left|b\right\rangle\}$ appears in the analysis as a description of the distinguishable effects of unitaries $\hat{U}(k)$ with eigenstates $\left|a\right\rangle$. The relation between the basis sets $\{\left|a\right\rangle\}$ and $\{\left|b\right\rangle\}$ thus originates from the relation between phase changes and experimentally observable changes in quantum mechanics. It might be worth noting that the same fundamental relation between phases and observable output is expressed by the quantum Fourier transform that appears in Shor’s algorithm [19]. In general, the quantum formalism relates information and dynamics by describing quantum processes as Hilbert space operators, where the same algebra is used for the representation of information gain by projection operators and for the representation of disturbance by unitary operators. The fundamental relation between information and disturbance formulated in this work thus captures a fundamental aspect of quantum mechanics with no correspondence in either classical physics or classical information theory.
Acknowledgments
This work was supported by ERATO, Japan Science and Technology Agency (JPMJER2402).